OpenStudy (anonymous):

Use a common denominator to add or subtract. Simplify your answer. : 3/x - 4/x^2 - 9x

OpenStudy (anonymous):

is this 3 4 --- - ---- - 9x x x^2

OpenStudy (anonymous):

(3/x - 4/x^2 - 9x)*x^2 = 3x - 4 -9x^3

OpenStudy (anonymous):

If my post is correct then take suzi20's answer over x^2 and you will have your answer

OpenStudy (anonymous):

no 4 ------- x^2 - 9x

OpenStudy (anonymous):

oh.. then you find the common denominator x = x and x^2 - 9x = x(x - 9) so your common denominator is x(x - 9) (x-9)3 4 ------- - ------- x(x-9) x(x-9) 3x - 27 4 ------ - -------- x(x - 9) x(x-9) 3x - 31 ------- x^2 - 9x

OpenStudy (anonymous):

3(x^2-9x) 4x -------- - -------- x(x^2-9x) x(x^2-9x)

OpenStudy (anonymous):

oh i'm late hahhaa

OpenStudy (anonymous):

thanks so much :D